NEED SOLUTION STEP BY STEP PLZ

Peter has an endowment of 3 units of good x and 5 units of good y. He can buy and sell x at a price of $100 and y at a price of $200. He receives an income of $700 as alimony from a former spouse.

a. Calculate the amount of x that he could afford if he bought only x and the amount of y he could afford if he bought only y.

b. Write an equation for Peter’s budget.
Answer #1

Endowment is 3 units of good x and 5 units of good y.

Price of good x= $100

Price of good y = $200

So, from this Peter's income = $100(3) + $200(5)= $1300

And he receives $700 as alimony from a former spouse.

So, now Peter has total income of $(1300+700)= $2000

(a) If Peter bought only good x , he could afford= $2000/$100 = 20 units.

And if Peter bought only good y ,he could afford = $2000/$200= 10 units.

(b) Peter's budget equation : Px X + Py Y = I

100 X + 200 Y = 2000